OpenStudy (anonymous):

2. The equation gives the distance, D, in miles that a person can see to the horizon from a height, h, in feet. a. Solve this equation for h. b. Long’s Peak in Rocky Mountain National Park, is 14,255 feet in elevation. How far can you see to the horizon from the top of Long’s Peak? Can you see Cheyenne, Wyoming (about 89 miles away)? Explain your answer.

OpenStudy (saifoo.khan):

did u tried to solve?

OpenStudy (saifoo.khan):

please dont make me give u direct answers. LOL.

OpenStudy (anonymous):

Yes would it be 14,255/89=160.17

OpenStudy (anonymous):

Do I divide

OpenStudy (anonymous):

I really feel stupid like I am not doing these right lol

OpenStudy (saifoo.khan):

surely u arent.

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

LOL

OpenStudy (saifoo.khan):

do the a) part first.

OpenStudy (anonymous):

k

OpenStudy (saifoo.khan):

can u make the equation?

OpenStudy (anonymous):

D=0.3\[\sqrt{14255}\]

OpenStudy (anonymous):

\[D=0.3\sqrt{14255}\]

OpenStudy (anonymous):

that is it give me a sec

OpenStudy (anonymous):

D=119.39

OpenStudy (anonymous):

is that close but how do I factor in the 0.3

OpenStudy (saifoo.khan):

u r going the wrong way, sorry for that.

OpenStudy (saifoo.khan):

Always try to solve the 1st part 1st. u have to make an equation first.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2. The equation gives the distance, D, in miles that a person can see to the horizon from a height, h, in feet. a. Solve this equation for h.

OpenStudy (saifoo.khan):

Yea, make a equation first.

OpenStudy (saifoo.khan):

like u did in the last ques, for a part.

OpenStudy (anonymous):

\[Here is the equation D=0.3\sqrt{14,255}\]

OpenStudy (anonymous):

How would I solve this

OpenStudy (anonymous):

using a calculater or by hand

OpenStudy (saifoo.khan):

Ok, im giving u the answer for the first part, the next one u will do by youself... DEAL?

OpenStudy (anonymous):

could you just explain to me how I get the answer then I will try it and you tell me if I did it right ?

OpenStudy (anonymous):

cause I want to learn

OpenStudy (saifoo.khan):

That's great. Ok.

OpenStudy (saifoo.khan):

Check your doc file, u missed one thing.

OpenStudy (anonymous):

2. The equation gives the distance, D, in miles that a person can see to the horizon from a height, h, in feet. a. Solve this equation for h. b. Long’s Peak in Rocky Mountain National Park, is 14,255 feet in elevation. How far can you see to the horizon from the top of Long’s Peak? Can you see Cheyenne, Wyoming (about 89 miles away)? Explain your answer.

OpenStudy (anonymous):

This is all it gives me

OpenStudy (saifoo.khan):

STILL NO.

OpenStudy (anonymous):

and the equation I put above that has 0.3 in it

OpenStudy (anonymous):

\[D=0.3\sqrt{h}\]

OpenStudy (saifoo.khan):

\[D = 1.2\sqrt{h}\] dont u have this?

OpenStudy (anonymous):

nope it is using 0.3

OpenStudy (anonymous):

not 1.2

OpenStudy (saifoo.khan):

Alright, so look now, the solution i gave u for the last one are not for that, that is a similar one, there is a bit difference.. in numericals. this one is also different one.

OpenStudy (anonymous):

k

OpenStudy (saifoo.khan):

now let me LOOK.

OpenStudy (anonymous):

k

OpenStudy (saifoo.khan):

u have to solve for h for the first part.q

OpenStudy (saifoo.khan):

try solving.

OpenStudy (saifoo.khan):

did u find the way out? (sorry for confusing u)

OpenStudy (anonymous):

q?

OpenStudy (saifoo.khan):

typo.

OpenStudy (anonymous):

oh ok

OpenStudy (saifoo.khan):

idk what is happening me today! lol

OpenStudy (anonymous):

lol I got D= 47516.67 not sure if I did this right but I took 0.3 divided into 14255 to get the answer

OpenStudy (saifoo.khan):

\[\sqrt{h}=\frac{D}{0.3}\] take a square root on both sides. \[h = \frac{D^2}{0.09}\]

OpenStudy (anonymous):

0.00.81

OpenStudy (anonymous):

0.0081

OpenStudy (saifoo.khan):

No. =0.3^2 =0.09

OpenStudy (anonymous):

oh lol ok got it now then I take 0.09 and divide into 14255?

OpenStudy (saifoo.khan):

now, u need to find the value for D.

OpenStudy (anonymous):

do I take 0.09 divided into 14255 which is h to get that?

OpenStudy (saifoo.khan):

No, u have to multiply.

OpenStudy (anonymous):

0.09*14255=1282.95

OpenStudy (anonymous):

So D=1282.95

OpenStudy (saifoo.khan):

That is D^2 = 1282.95

OpenStudy (anonymous):

so now I have to take that answer to the second power ?

OpenStudy (saifoo.khan):

take a root on both sides. this will cancel the square of D and it will go like: D = sqrt(1282.95)

OpenStudy (anonymous):

35.812 did I do that right

OpenStudy (saifoo.khan):

35.818

OpenStudy (anonymous):

yeah I did get 35.818 to I just redid it on the calculater sorry

OpenStudy (saifoo.khan):

No, prob.. but u r right! :)

OpenStudy (anonymous):

sweet ty

OpenStudy (anonymous):

2355805350*3962.946591ˉ²=15704945.6831185

OpenStudy (anonymous):

did I get the correct answer for this one to

OpenStudy (saifoo.khan):

what is that??

OpenStudy (anonymous):

LOL sorry to confuse you I just wanna know if I did the math right

OpenStudy (saifoo.khan):

haha!